# Neuro-Symbolic Financial Reasoning via Deterministic Fact Ledgers and Adversarial Low-Latency Hallucination Detector

### 04 Mar 2026 v1

@@ Abstract
Standard Retrieval-Augmented Generation (RAG) architectures fail in high-stakes financial domains due to two fundamental limitations: the inherent arithmetic incompetence of Large Language Models (LLMs) and the distributional semantic conflation of dense vector retrieval (e.g., mapping “Net Income” to “Net Sales” due to contextual proximity). In deterministic domains, a 99% accuracy rate yields 0% operational trust. To achieve zero-hallucination financial reasoning, we introduce the Verifiable Numerical Reasoning Agent (VeNRA). VeNRA shifts the RAG paradigm from retrieving probabilistic text to retrieving deterministic variables via a strictly typed Universal Fact Ledger (UFL), mathematically bounded by a novel Double-Lock Grounding algorithm. Recognizing that upstream parsing anomalies inevitably occur, we introduce the VeNRA Sentinel: a 3-billion parameter SLM trained to forensically audit Python execution traces with only one token test budget. To train this model, we avoid traditional generative hallucination datasets in favor of Adversarial Simulation, programmatically sabotaging golden financial records to simulate production-level “Ecological Errors” (e.g., Logic Code Lies and Numeric Neighbor Traps). Finally, to optimize the Sentinel under strict latency budgets, we utilize a single-pass classification paradigm with optional post thinking for debug. We identify the phenomenon of Loss Dilution in Reverse-Chain-of-Thought training and present a novel, OOM-safe Micro-Chunking loss algorithm to stabilize gradients under extreme differential penalization.
@@ 1 Introduction
Figure 1: The VeNRA Neuro-Symbolic Paradigm. (Top) Standard RAG suffers from “Stochastic Inaccuracy” due to probabilistic arithmetic and vector conflation. (Bottom) VeNRA parses PDFs into a Universal Fact Ledger (UFL). A Query Navigator retrieves specific variables via a Lexical Gate. An Architect LLM generates a deterministic Python trace ($T$). Crucially, the 33B Sentinel model forensically audits the logic trace before surfacing the answer.
@@
In this paper, we argue that reliable financial AI cannot be achieved purely through parametric scaling or generative evaluation. Instead, it requires Neuro-Symbolic Cognitive Offloading and Forensic Auditing. We introduce the VeNRA [^1] The code is available in Github: https://github.com/pagand/VeNRA (Verifiable Numerical Reasoning Agent) framework, an end-to-end ecosystem designed to prevent, simulate, and detect financial hallucinations. To mitigate hallucinations, VeNRA replaces unstructured text retrieval with the Universal Fact Ledger (UFL). By strictly extracting text into typed variables backed by Double-Lock Grounding (requiring both character-offset alignment and semantic schema validation), and gating semantic retrieval with deterministic lexical constraints, we relegate the LLM entirely to the role of a deterministic Python code architect. To detect slipped hallucinations, we introduce the VeNRA Sentinel, a locally deployable $3$-B parameter SLM capable of auditing complex mathematical traces in <50<50ms, zero-shot, zero test-time compute. To train the Sentinel, we pioneer Adversarial Simulation, rejecting generative noise in favor of VeNRA-Data: a 10,000-sample dataset constructed by programmatically sabotaging golden financial records to recreate the exact mechanical failures of production RAG systems.
@@
  • Neuro-Symbolic Architecture: The Universal Fact Ledger (UFL) with Double-Lock Grounding and Lexical Pre-Filtering, virtually eliminating extraction hallucination and distributional vector conflation.
@@
  • Forensic Auditing Systems Engineering [^3] https://huggingface.co/spaces/pagand/VeNRA_halDet: The VeNRA Sentinel, a 3B parameter judge achieving low latency verification. We contribute the OOM-safe Micro-Chunking Trainer, mathematically solving the gradient dilution problem in Reverse-CoT architectures.
@@ 2.2 Numerical Reasoning and Neuro-Symbolic Agents
However, benchmarks like FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021) reveal that even code-generating models struggle with grounding; frequently executing correct logic on hallucinated input variables. VeNRA extends the PAL paradigm by enforcing a stricter separation of concerns: the LLM acts solely as a Code Architect operating over a strictly typed Universal Fact Ledger (UFL), preventing the injection of unsupported scalars into the computation graph.
@@ 3 The VeNRA Architecture: Grounded Neuro-Symbolic RAG
To mitigate these fundamental flaws, we introduce VeNRA to shift the RAG paradigm from “Retrieving Text” to “Retrieving Variables”. Rather than forcing a single LLM to simultaneously act as a semantic search engine, an unstructured parser, and an arithmetic calculator, VeNRA strictly decouples these cognitive burdens. Unstructured documents are deterministically parsed into a strongly typed Universal Fact Ledger (UFL), mathematically grounded, and subsequently queried by an LLM constrained strictly to Python code generation.
3.1 Structure-Preserving Ingestion and the Universal Fact Ledger
The Universal Fact Ledger (UFL) serves as the foundational data contract for the VeNRA ecosystem, acting simultaneously as a computational table for arithmetic and a relational graph for semantic traversal. To populate the UFL from unstructured 10-K filings without succumbing to “Context Rot” Hong et al. (2025), the loss of structural lineage and coreference, we employ a deterministic ingestion pipeline.
Trailing-Buffer Semantic Chunking. Standard recursive chunking mechanisms frequently sever cross-boundary coreferences (e.g., a table row referencing “See Note 12”, where the note text resides in the subsequent chunk). To preserve localized context, we developed a newline-preference trailing-buffer algorithm. Let $C_{i}$ be the $i$-th chunk of text, bounded by a character threshold $\tau=3000$. Rather than splitting at an arbitrary character index, the algorithm splits at the most recent newline character. For every generated text block $C_{i}$, a trailing buffer $b_{i-1}$ consisting of the final $k=300$ characters of the preceding block $C_{i-1}$ is injected as a prefix:
@@ 3.1 Structure-Preserving Ingestion and the Universal Fact Ledger
Deterministic Table Melting. Generative LLMs exhibit severe unreliability when extracting 2D grid data Zhu et al. (2021). Consequently, VeNRA bypasses LLM extraction for tables entirely, employing deterministic Pandas routines. Raw markdown tables are parsed using a hierarchical disambiguation stack. Indentation markers (e.g., &nbsp;) are computed to generate fully qualified schema paths (e.g., Assets > Current Assets > Inventory), preventing the collapse of distinct sub-categories into ambiguous metrics like “Other.” Furthermore, row-level overrides dynamically intercept global column scale factors; metrics containing substrings such as “per share” enforce a $1.0$ multiplier and a unit classification of USD/Share, ensuring mathematical invariants are preserved.
@@ 3.4 Hybrid Lexical-Semantic Retrieval
Once the Universal Fact Ledger (UFL) is populated, the system must retrieve the precise variables required to satisfy a user query. Pure semantic vector search is fundamentally ill-suited for financial retrieval because dense embedding models (e.g., bge-m3) rely on distributional semantics. Consequently, they frequently group mathematically opposite terms by contextual proximity. A query for “Net Sales” will consistently retrieve chunks discussing “Net Income” with high cosine similarity, leading to cascading arithmetic failures downstream.
To overcome this, VeNRA introduces a Hybrid Lexical-Semantic Retriever, governed by an LLM-based Query Navigator and a deterministic Lexical Gate.
Query Navigation and Schema Mapping. Upon receiving a query, a Navigator SLM evaluates the user intent against a dynamically injected schema summary, which maps canonical entity IDs and high-frequency metric names. The Navigator outputs a bipartite RetrievalPlan containing: (1) a structured UFL filter (e.g., Entity ID, Year, Metric Keywords, and Nuance conditions), and (2) a “Vector Hypothesis”, an idealized text snippet theoretically containing the answer.
Lexical Pre-Filtering (The Semantic Gate). To prevent distributional semantic conflation, all vector retrieval candidates (and fuzzy UFL metric matches) must pass a deterministic Lexical Gate before entering the context window.
@@ 3.5 Program-Aided Execution and Context Assembly
Two-Pass Deterministic Orchestration. VeNRA executes reasoning via a dual-pass neuro-symbolic flow:
@@
Runtime Fallback Protocol: “Trust Text Over Data”. Despite rigorous ingestion grounding, optical character recognition (OCR), failure in UFL underlying LLM, and PDF parsing anomalies inevitably occur. VeNRA implements a programmatic fault-tolerance mechanism. The Architect agent is explicitly instructed via its system prompt to cross-reference the structured UFL rows against the provided raw text chunks. If the agent detects that the UFL data is malformed (e.g., an unaligned temporal column) or mathematically contradicts the raw text, it programmatically discards the UFL variable. Instead, the Python code is written to extract and compute the values directly from the provided source text string, guaranteeing system resilience.
@@ 4 VeNRA-Data: Adversarial Simulation of RAG Failures
To effectively train a low-latency Sentinel model to audit the outputs of the VeNRA architecture, we require a dataset that accurately reflects the failure modes of financial RAG systems. We introduce VeNRA-Data, a heavily curated, 10,000-sample dataset comprising contrastive pairs of grounded and ungrounded financial reasoning traces.
@@ 4.3 The Teacher-Auditor Protocol and Noise Injection
Dual-Path Cause-Effect Validation. A critical algorithmic hurdle in automated auditing is the decoupling of cause and effect. In a “Logic Code Lie” (Section 4.2), the Saboteur mutates an input variable ($v\leftarrow d$). However, the Teacher model frequently flags the output scalar (the final computed answer in $S^{\prime}$) as the error span, rather than the injected input $d$. A naive string-matching validation script falsely rejects these valid sabotages because the cause ($d$) does not equal the effect ($S^{\prime}$).
@@
    Effect Match: The fuzzy token set ratio between $d$ and $E_{teacher}$ exceeds a threshold $\tau_{fuzz}>60$.
@@
Axiomatic Noise Injection. To train the GENERAL class, we utilize TruthfulQA’s finance and economics splits. Initially, these records contain an empty context ($C=\emptyset$). However, training an autoregressive model on this representation induces a spurious correlation: the model learns the structural shortcut $C=\emptyset\implies\texttt{GENERAL}$, bypassing semantic reasoning entirely. To force true logic-awareness, we implement Axiomatic Noise Injection. For every general axiom, we sample a subset of completely irrelevant financial text chunks from a distractor pool ($C_{distractor}$). The model must learn to actively ignore the distractor context and rely on its parametric memory when the trace signals GENERAL_KNOWLEDGE, preventing structural overfitting.
The “Virtual Context” Protocol and HITL Refinement. For highly ambiguous samples where the automated validation yields low confidence, we utilize an AI Proxy Refiner. To minimize context-window bloat and token costs, the Refiner operates under a “Virtual Context” protocol: it is denied access to the raw context $C$. Instead, it relies entirely on the Teacher’s $\Theta_{teacher}$ block as its “eyes.” Guided by a strict meta-prompt, the Refiner surgically corrects nomenclature confusions, sign-convention inversions (e.g., treating a cash outflow as an inflow), and boilerplate evasions, effectively overruling the Teacher model when it exhibits “math-blindness.”. If the result was still ambigious, we tag for human annotation via the specific UI shown in Fig. 2.
@@ 5 VeNRA Sentinel: Forensic Auditing under Latency Constraints
5.1 The Sytem 1.5 Paradigm and Token Orthogonality
@@
To satisfy latency constrain, we inspired from SALSA (Single-pass Autoregressive LLM Structured Classification) framework, formatting the training targets as Reverse-CoT:
@@ 5.2 The Effects of Loss Dilution
To restore the gradient signal of the verdict, we must apply a large differential weight to the label token. However, implementing a weighted loss via PyTorch’s CrossEntropyLoss(reduction=’none’) on a $3$-B parameter model triggers immediate Out-Of-Memory (OOM) failures on a single GPU (with <50GB of VRAM) [^5] For a micro-batch size $B=2$ and sequence length $S=4096$, calculating the unreduced loss instantiates a tensor of size $B\times S\times|V|]$. At FP32 precision (required for loss stability), this single intermediate tensor consumes $\approx 1.25$ Billion floats ($\approx 5$ GB). Combined with gradients and optimizer states, this vastly exceeds VRAM capacity.. To solve this, we employed a custom Micro-Chunking Trainer. We bypass the model’s internal loss calculation, flattening the shift-logits and shift-labels. We then iteratively slice the sequence into micro-chunks of size $c=512$ entirely within the loss computation loop. For a flat sequence length $T=B\times S$, the weighted loss is accumulated safely:
@@ 5.3 Dynamic Loss Clamping and Differential Weighting
While the Micro-Chunking Trainer enables weighted loss, multiplying the label gradient by $50\times$ introduces severe topological instability (Gradient Seesaw). Imposing a large penalty on a confidently wrong prediction (where $P(\hat{y})\approx 0$) causes the cross-entropy loss to spike asymptotically [^6] During early training, we observed massive gradient norm explosions ($\|g\|>72,000$), causing the model to violently oscillate, updating from predicting $100\%$ Fake in one step to $100\%$ Found in the next.. To stabilize convergence, we introduced Dynamic Loss Clamping, explicitly capping the maximum penalization:

### 08 Mar 2026 v2

@@ Abstract
Standard Retrieval-Augmented Generation (RAG) architectures fail in high-stakes financial domains due to two fundamental limitations: the inherent arithmetic incompetence of Large Language Models (LLMs) and the distributional semantic conflation of dense vector retrieval (e.g., mapping Net Income to Net Sales due to contextual proximity). In deterministic domains, a $99\%$ accuracy rate yields $0\%$ operational trust. To achieve zero-hallucination financial reasoning, we introduce the Verifiable Numerical Reasoning Agent (VeNRA). VeNRA shifts the RAG paradigm from retrieving probabilistic text to retrieving deterministic variables via a strictly typed Universal Fact Ledger (UFL). We mathematically bound this ledger using a novel Double-Lock Grounding algorithm. Coupled with deterministic Python execution, this neuro-symbolic routing compresses systemic hallucination rates to a near-zero $1.2\%$. Recognising that upstream parsing anomalies inevitably occur, we introduce the VeNRA Sentinel: a 3-billion parameter SLM trained to forensically audit candidate using a single-token inference budget with optional post-hoc reasoning. To train the Sentinel, we steer away from traditional hallucination datasets in favour of Adversarial Simulation, programmatically sabotaging financial records to simulate Ecological Errors. The compact Sentinel consequently outperforms 7070B++ frontier models in error detection. Through Loss Dilution phenomenon in Reverse-CoT training, we present a novel Micro-Chunking loss algorithm to stabilise gradients under extreme verdict penalisation, yielding a $28{\times}$ latency speedup without sacrificing forensic rigor.
@@ 1 Introduction
Figure 1: The VeNRA Neuro-Symbolic Paradigm. (Top) Standard RAG suffers from “Stochastic Inaccuracy” due to probabilistic arithmetic and vector conflation. (Bottom) VeNRA parses PDFs into a Universal Fact Ledger (UFL). A Query Navigator retrieves specific variables via a Lexical Gate. An Architect LLM generates a deterministic Python trace. Crucially, the 33B Sentinel model forensically audits the logic trace before surfacing the answer.
@@
In this paper, we argue that reliable financial AI cannot be achieved purely through parametric scaling or generative evaluation. Instead, it requires Neuro-Symbolic Cognitive Offloading and Forensic Auditing. We introduce the VeNRA [^1] The code is available at Github: https://github.com/pagand/VeNRA (Verifiable Numerical Reasoning Agent) framework, an end-to-end ecosystem designed to prevent, simulate, and detect financial hallucinations. To mitigate hallucinations, VeNRA replaces unstructured text retrieval with the UFL. By strictly extracting text into typed variables backed by Double-Lock Grounding (requiring both character-offset alignment and semantic schema validation), and gating semantic retrieval with deterministic lexical constraints, we relegate the LLM entirely to the role of a deterministic Python code architect. To detect slipped hallucinations, we introduce the VeNRA Sentinel, a locally deployable $3$-B parameter SLM capable of auditing complex mathematical traces in <50<50ms, zero-shot, zero test-time compute. To train the Sentinel, we pioneer Adversarial Simulation, rejecting generative noise in favor of VeNRA-Data: a 10,000-sample dataset constructed by programmatically sabotaging golden financial records to recreate the exact mechanical failures of production RAG systems.
@@
  • Neuro-Symbolic Architecture: The UFL with Double-Lock Grounding and Lexical Pre-Filtering, virtually eliminating extraction hallucination and distributional vector conflation.
@@
  • Forensic Auditing Systems Engineering [^3] https://huggingface.co/spaces/pagand/VeNRA_halDet: The VeNRA Sentinel, a 3B parameter judge achieving low latency verification. We contribute the Micro-Chunking Trainer, mathematically solving the gradient dilution problem in Reverse-CoT architectures.
@@ 2.2 Numerical Reasoning and Neuro-Symbolic Agents
However, benchmarks like FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021) reveal that even code-generating models struggle with grounding; frequently executing correct logic on hallucinated input variables. VeNRA extends the PAL paradigm by enforcing a stricter separation of concerns: the LLM acts solely as a Code Architect operating over a strictly typed UFL, preventing the injection of unsupported scalars into the computation graph.
@@ 3 The VeNRA Architecture: Grounded Neuro-Symbolic RAG
To mitigate these fundamental flaws, we introduce VeNRA to shift the RAG paradigm from “Retrieving Text” to “Retrieving Variables”. Rather than forcing a single LLM to simultaneously act as a semantic search engine, an unstructured parser, and an arithmetic calculator, VeNRA strictly decouples these cognitive burdens. Unstructured documents are deterministically parsed into a strongly typed UFL, mathematically grounded, and subsequently queried by an LLM constrained strictly to Python code generation.
3.1 Ingestion via Universal Fact Ledger
The UFL serves as the foundational data contract for the VeNRA ecosystem, acting simultaneously as a computational table for arithmetic and a relational graph for semantic traversal. To populate the UFL from unstructured 10-K filings without succumbing to “Context Rot” Hong et al. (2025), the loss of structural lineage and coreference, we employ a deterministic ingestion pipeline.
Standard recursive chunking mechanisms frequently sever cross-boundary coreferences (e.g., a table row referencing “See Note 12”, where the note text resides in the subsequent chunk). To preserve localized context, we developed a newline-preference trailing-buffer algorithm. Let $C_{i}$ be the $i$-th chunk of text, bounded by a character threshold $\tau=3000$. Rather than splitting at an arbitrary character index, the algorithm splits at the most recent newline character. For every generated text block $C_{i}$, a trailing buffer $b_{i-1}$ consisting of the final $k=300$ characters of the preceding block $C_{i-1}$ is injected as a prefix:
@@ 3.1 Ingestion via Universal Fact Ledger
Generative LLMs exhibit severe unreliability when extracting 2D grid data Zhu et al. (2021). Consequently, VeNRA bypasses LLM extraction for tables entirely, employing deterministic Pandas routines. Raw markdown tables are parsed using a hierarchical disambiguation stack. Indentation markers (e.g., &nbsp;) are computed to generate fully qualified schema paths (e.g., Assets > Current Assets > Inventory), preventing the collapse of distinct sub-categories into ambiguous metrics like “Other.” Furthermore, row-level overrides dynamically intercept global column scale factors; metrics containing substrings such as “per share” enforce a $1.0$ multiplier and a unit classification of USD/Share, ensuring mathematical invariants are preserved.
@@ 3.4 Hybrid Lexical-Semantic Retrieval
Once the UFL is populated, the system must retrieve the precise variables required to satisfy a user query. Pure semantic vector search is fundamentally ill-suited for financial retrieval because dense embedding models (e.g., bge-m3) rely on distributional semantics. Consequently, they frequently group mathematically opposite terms by contextual proximity. A query for “Net Sales” will consistently retrieve chunks discussing “Net Income” with high cosine similarity, leading to cascading arithmetic failures downstream.
To overcome this, VeNRA introduces a Hybrid Lexical-Semantic Retriever, governed by an LLM-based Query Navigator Agand et al. (2024) and a deterministic Lexical Gate.
Upon receiving a query, a Navigator SLM evaluates the user intent against a dynamically injected schema summary, which maps canonical entity IDs and high-frequency metric names. The Navigator outputs a bipartite RetrievalPlan containing: (1) a structured UFL filter (e.g., Entity ID, Year, Metric Keywords, and Nuance conditions), and (2) a “Vector Hypothesis”, an idealized text snippet theoretically containing the answer.
To prevent distributional semantic conflation, all vector retrieval candidates (and fuzzy UFL metric matches) must pass a deterministic Lexical Gate before entering the context window.
@@ 3.5 Program-Aided Execution and Context Assembly
VeNRA executes reasoning via a dual-pass neuro-symbolic flow:
@@
Despite rigorous ingestion grounding, optical character recognition (OCR), failure in UFL underlying LLM, and PDF parsing anomalies inevitably occur. VeNRA implements a programmatic fault-tolerance mechanism. The Architect agent is explicitly instructed via its system prompt to cross-reference the structured UFL rows against the provided raw text chunks. If the agent detects that the UFL data is malformed (e.g., an unaligned temporal column) or mathematically contradicts the raw text, it programmatically discards the UFL variable. Instead, the Python code is written to extract and compute the values directly from the provided source text string, guaranteeing system resilience.
@@ 4 VeNRA-Data: Adversarial Simulation of RAG Failures
To effectively train a low-latency Sentinel model to audit the outputs of the VeNRA architecture, we require a dataset that accurately reflects the failure modes of financial RAG systems. We introduce VeNRA-Data, a heavily curated dataset comprising contrastive pairs of grounded and ungrounded financial reasoning traces.
@@ 4.3 The Teacher-Auditor Protocol and Noise Injection
A critical algorithmic hurdle in automated auditing is the decoupling of cause and effect. In a “Logic Code Lie” (Section 4.2), the Saboteur mutates an input variable ($v\leftarrow d$). However, the Teacher model frequently flags the output scalar (the final computed answer in $S^{\prime}$) as the error span, rather than the injected input $d$. A naive string-matching validation script falsely rejects these valid sabotages because the cause ($d$) does not equal the effect ($S^{\prime}$).
@@
    Effect Match: The fuzzy token set ratio between $d$ and $E_{teacher}$ exceeds a threshold.
@@
To train the GENERAL class, we utilize TruthfulQA’s finance and economics splits. Initially, these records contain an empty context ($C=\emptyset$). However, training an autoregressive model on this representation induces a spurious correlation: the model learns the structural shortcut $C=\emptyset\implies\texttt{GENERAL}$, bypassing semantic reasoning entirely. To force true logic-awareness, we implement Axiomatic Noise Injection. For every general axiom, we sample a subset of completely irrelevant financial text chunks from a distractor pool ($C_{distractor}$). The model must learn to actively ignore the distractor context and rely on its parametric memory when the trace signals GENERAL_KNOWLEDGE, preventing structural overfitting. For highly ambiguous samples where the automated validation yields low confidence, we utilize an AI Proxy Refiner (More on Appendix.B).
@@ 5 VeNRA Sentinel: Forensic Auditing under Latency Constraints
5.1 The System 1.5 Paradigm and Token Orthogonality
@@
To satisfy latency constrain, we inspired from SALSA (Single-pass Autoregressive LLM Structured Classification) Berdichevsky et al. (2025) framework, formatting the training targets as Reverse-CoT:
@@ 5.2 The Effects of Loss Dilution
To restore the gradient signal of the verdict, we must apply a large differential weight to the label token. However, implementing a weighted loss via PyTorch’s CrossEntropyLoss(reduction=’none’) on a $3$-B parameter model triggers immediate OOM failures on a single GPU (with <50GB of VRAM) [^5] For a micro-batch size $B=2$ and sequence length $S=4096$, calculating the unreduced loss instantiates a tensor of size $B\times S\times|V|]$. At FP32 precision (required for loss stability), this single intermediate tensor consumes $\approx 1.25$ Billion floats ($\approx 5$ GB). Combined with gradients and optimizer states, this vastly exceeds VRAM capacity.. To solve this, we employed a custom Micro-Chunking Trainer. We bypass the model’s internal loss calculation, flattening the shift-logits and shift-labels. We then iteratively slice the sequence into micro-chunks of size $c=512$ entirely within the loss computation loop. For a flat sequence length $T=B\times S$, the weighted loss is accumulated safely:
@@ 5.3 Dynamic Loss Clamping and Differential Weighting
While the Micro-Chunking Trainer enables weighted loss, multiplying the label gradient by $50\times$ introduces severe topological instability (Gradient Seesaw). Imposing a large penalty on a confidently wrong prediction (where $P(\hat{y})\approx 0$) causes the cross-entropy loss to spike asymptotically [^6] During early training, we observed massive gradient norm explosions, causing the model to violently oscillate, updating from predicting $100\%$ Fake in one step to $100\%$ Found in the next.. To stabilize convergence, we introduced Dynamic Loss Clamping, explicitly capping the maximum penalization: